# S2S-Net: Addressing the Domain Gap of Heterogeneous Sensor Systems in LiDAR-Based Collective Perception

### v1

@@ VI Conclusion and Future Work
In this work, we presented S2S-Net, a sensor-domain robust fusion architecture using sparse voxel grids as a compact and unified environment representation for collective perception. This is the first work addressing the Sensor2Sensor domain gap in V2V collective perception.
We demonstrated that S2S-Net is very robust to changes in the sensor domain and that the local perception could be significantly improved for all classes no matter which sensor data was shared with the ego vehicle. Even when changing the testing domain between very dissimilar sensors or selecting different sensors for each CAV, no significant drop in perception performance could be observed that could be traced back to the sensor domain gap.
For future research, we will extend our evaluation to different voxel grid resolutions. Furthermore, we want to incorporate other fusion methods into our study to further characterize the sensor domain gap in V2V collective perception. Moreover, we aim to conduct experiments incorporation infrastructural sensors to also study the domain gap arising from different placements of sensors.

### v2

@@ VI Conclusion and Future Work
In this work, we presented S2S-Net, a sensor-domain robust fusion architecture using sparse voxel grids as a compact and unified environment representation for collective perception. This is the first work addressing the Sensor2Sensor domain gap in V2V collective perception. This study showed, that all the evaluated state-of-the-art methods highly suffer from the Sensor2Sensor domain gap. We demonstrated that S2S-Net is very robust to changes in the sensor domain and that it could significantly outperform all other evaluated methods. Especially when the sensor domains were highly different, S2S-Net achieved up to 44 percentage points higher average precision than the other methods. For future research, we will extend our evaluation to different voxel grid resolutions. Furthermore, we want to incorporate more fusion methods into our study to further characterize the sensor domain gap in V2V collective perception. Moreover, we aim to conduct experiments incorporation infrastructural sensors to also study the domain gap arising from different placements of sensors.